Two numbers have a product of 80. Their sum is 21. What are the numbers?
Answer by checkley71(8403)

X+Y=21 OR X=21-Y
X*Y=80
(21-Y)Y=80
21Y-Y^2-80=0
-Y^2+21Y-80=0
Y^2-21+80=0
(Y-15)(Y-6)=0
Y-15=0
Y=15 ANSWER.
Y-6=0
Y=6 ANSWER.
15+6=21
15*6=80
80=80